OpenStudy (anonymous):

solve for x and y 5x-4y=43 2x+3y=8 explain plz.

OpenStudy (anonymous):

you can sum the two equations, term by term, obtaining 7x-y=51 y=7x-51 and substitute y in one of them e.g. the second one 2x+21x-153=8 Then you solve for x 23x=161 x=7 and substitute it in the other (the first) 35-4y=43 y=-8/4 y=2

OpenStudy (anonymous):

and how r the equations summed? hm???

OpenStudy (anonymous):

term by term in our case: 5x-4y +2x+3y = 43 +8

OpenStudy (anonymous):

You can also subtract one from another. You always get a true equation